Circular No. 9013

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2009 B1 (BOATTINI) A. Boattini reports his discovery of a comet with a strong condensa-

tion, a 10′′ coma, and a wide tail ∼ 30′′ long in p.a. 60^{o} on images taken with the Catalina Sky Survey’s 0.68-m Schmidt telescope (discovery observation tabulated below). Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, numerous other CCD astrometrists have commented on the object’s cometary appearance. W. H. Ryan (Magdalena Ridge Observatory, Socorro, NM, U.S.A., 2.4-m f/8.9 reflector, Jan. 21.2 UT) reports a “fuzzy appearance” with a faint tail in p.a. ≈ 60^{o}. J. Bedient (2.0-m “Faulkes Telescope North” at Haleakala, Jan. 21.3) writes that a co-added 180-s R-band exposure shows a 6′′ coma and a 10′′-long tail in p.a. 55^{o}. Five stacked 300-s exposures taken by T. Kryachko and S. Korotkiy (Astrotel Observatory, Kazan State University, 0.30-m Takahashi telescope, Jan. 21.7) show a faint 18′′ tail in p.a. 60^{o} and a coma diameter of ∼ 9′′. L. Buzzi (Varese, Italy, 0.60-m reflector, Jan. 21.8) finds the comet to be diffuse with an asymmetric coma ∼ 10′′ wide with a central condensation but no noticeable tail. R. Holmes (Charleston, IL, U.S.A., 0.81-m astrograph, Jan. 22.0) reports that a 15-min stack of images reveals a nuclear condensation 9′′.0 wide and a broad, diffuse, 21′′.8 tail in p.a. 78^{o}.9. D. T. Durig and M. F. Mansfield (Sewanee, TN, U.S.A., 0.30-m reflector, Jan. 22.1) notes that twenty stacked 120-s frames show a coma of diameter 8′′– 10′′ with a hint of an 10′′ tail in p.a. 75^{o}–80^{o}. E. Guido, Castellammare di Stabia, Italy, writes that eighteen co-added 120-s unfiltered exposures that he took remotely with G. Sostero and P. Camilleri using a 0.25-m reflector near Mayhill, NM, U.S.A., on Jan. 22.1 show a coma diameter of ∼ 11′′ and a tail ∼ 15′′ long in p.a. 70^{o}.

2009 UT  α_{2000}  δ_{2000}  Mag. Observer
Jan. 21.07424 1^{h}48^{m}22.^{s}96 +34^{o}00′22′′.5 18.2 Boattini

The available astrometry, the following very preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2009-B46. It is possible that this comet is of short period.

T = 2009 Apr. 14.471 TT  ω = 158.^{o}062
Ω = 304.744 }2000.0
q = 1.96025 AU  i = 20.617
2009 January 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green